# Kenyon’s identities for the height function and compactified free field in the dimer model

### v1

@@ Abstract.
In a seminal paper published in 2000 Kenyon developed a method to study the height function of the planar dimer model via discrete complex analysis tools. The core of this method is a set of identities representing height correlations through the inverse Kasteleyn operator. Scaling limits of these identities (if exist) produce a set of correlation functions written in terms of a Dirac Green’s kernel with unknown boundary conditions. It was proven in [CLR1] that, under natural assumptions, these correlations always define a Gaussian free field in a simply connected domain. This was generalized to doubly connected domains in the recent work [ChelkakDeiman], where the field is shown to be a sum of Gaussian free field and a discrete Gaussian component. We generalize this result further to arbitrary bordered Riemann surfaces.
@@
A University of Helsinki.
@@ 1. Introduction
In his pioneering works [KenyonConfInvOfDominoTilings, KenyonGFF] Kenyon considered the dimer model sampled in a sequence of Temperleyan polygons $\Omega_{k}\subset\frac{1}{k}{\mathbb{Z}}^{2}$ that approximate a planar domain $\Omega\subset{\mathbb{C}}$ and proved the following two results. Assuming simply connectedness of $\Omega$, the centered dimer height functions converge to the (properly scaled) Gaussian free field with Dirichlet boundary conditions [KenyonGFF]. For multiply connected $\Omega$ the height gaps between boundary components of $\Omega_{k}$ converge to unknown conformally invariant quantities [KenyonConfInvOfDominoTilings].
@@
Once one is able to make this line of arguments rigorous, one still has to identify the right-hand side of (1.2). This can be done, for example, by identifying the boundary conditions of the Dirac operator inherited by the combinatorics of discrete domains along their boundaries. For instance, Temperleyan combinatorics considered in [KenyonConfInvOfDominoTilings, KenyonGFF] imposes Dirichlet boundary conditions. These boundary conditions make $f^{[\pm,\pm]}$ conformally covariant, which implies the conformal invariance of the correlations (1.2) and allows one to identify $\overline{h}$ with the Gaussian free field in the case of a simply connected $\Omega$. As shown in a recent work [nicoletti2025temperleyan], if $\Omega$ is multiply connected, these boundary conditions also allow one to express $f^{[\pm,\pm]}$ explicitly via the classical Szegö kernel on the Schottky double of $\Omega$ [Fay, Chapter VI]. This allows to fully determine the field $\overline{h}$ and, in particular, to identify the distribution of boundary height gaps considered by Kenyon in [KenyonConfInvOfDominoTilings].
Other examples of explicit boundary conditions were studied in [russkikh2018dimers, russkikh2020dominos]. In general, however, the boundary conditions of the dimer models in question may be quite nasty and do not produce conformally covariant scaling limits $f^{[\pm,\pm]}$, even if the correlations (1.2) are expected to be conformally invariant. Moreover, regardless the boundary conditions of the Dirac operator, the random field $\overline{h}$ with correlations (1.2) is expected to always belong to a certain universal family of random distributions corresponding to compactified free fields in $\Omega$. This poses a natural question about properties of $f^{[\pm,\pm]}$ that are sufficient to guarantee the conformal invariance of the right-hand side of (1.2) and, moreover, to identify the field $\overline{h}$ with a canonical object such as compactified free field.
@@
In the case of a simply connected domain a compelling answer to this question is given by [CLR1, Lemma 7.3] that asserts the following. Denote the right-hand side of (1.2) by $U_{m}$. By definition $U_{m}$ is a harmonic differential in each of the variables. Assume that $U_{2}$ and $U_{3}$ have Dirichlet boundary conditions (which one expects as $\overline{h}$ is supposed to vanish outside of the domain) and correct singularities in the bulk: that is, $U_{2}$ has a quadratic pole with a prescribed residue along the diagonal and $U_{3}$ is regular. Then there exists a gauge relating $f^{[\pm,\pm]}$ to the components of the Dirac Green’s kernel satisfying “spinor-type” boundary conditions, and, in particular, $U_{m}$’s coincide with multi-point correlations of the gradient of the Gaussian free field.
@@
When $\Omega$ is multiply connected with boundary components $B_{0},B_{1},\dots,B_{n},\ n\geq 1,$ well-known heuristics [gorin2021lectures, Lecture 24] suggest that the limit field $\overline{h}$ should be of the form
@@
Thus, generalizing [CLR1, Lemma 7.3] to multiply connected domains requires to determine the distribution of the discrete component $c$ additionally to the GFF component $\phi$. In the case of a doubly connected domain such generalization was very recently obtained in [ChelkakDeiman] using properties of the Weierstrass $\wp$-function. This, in particular, allowed the authors of [ChelkakDeiman] to study the dimer model on a cylinder with Temperleyan conditions of different ‘colors’ on two boundaries, in which case the Dirac Green’s kernel $f^{[\pm,\pm]}$ is not conformally covariant (see Section 5.2).
The goal of this note is to generalize this approach further: from simply and doubly connected setups to arbitrary multiply connected domains and bordered Riemann surfaces. Namely, we show that the algebraic structure of Kenyon’s identities (1.2) allows one to identify the form of the joint distribution of dimer height gaps between boundary components and thus provide a strong supporting evidence for [gorin2021lectures, Conjecture 24.1].
@@ 1.1. Main result: multiply connected domains
To formulate our first theorem we need several definitions. Let $U_{m}$ denote the right-hand side of (1.2) as before. Let us say that $h=\phi+\sum_{j=1}^{n}c_{j}\mathrm{hm}_{j}$ is a compactified free field with the shift $c_{0}\in{\mathbb{R}}^{n}$ if $\phi$ and $c$ are as described above. Recall that the right-hand side of (1.3) is $h-\operatorname{{\mathbb{E}}}h$. Given an integer vector $\mathfrak{m}\in{\mathbb{Z}}^{n}$ define
@@
For various applications it is also convenient to allow functions $f^{[\pm,\pm]}$ to have multiplicative monodromies. (For instance, this includes the setup of [ChelkakDeiman] in which case $f^{[\pm,\pm]}$ have monodromy 1-1.) Let us say that $f^{[\pm,\pm]}(z_{1},z_{2})$ are multivalued with multiplicative monodromy if $f^{[\pm,\pm]}(z_{1},z_{2})$ are functions defined for $(z_{1},z_{2})$ from the universal cover of $\Omega\times\Omega$, and there exist characters $\chi_{1,2}:\pi_{1}(\Omega)\to{\mathbb{T}}$ such that for every deck transformation $(\gamma_{1},\gamma_{2})\in\pi_{1}(\Omega)\times\pi_{1}(\Omega)$
(1.6) $(\gamma_{1},\gamma_{2})^{\ast}f^{[\pm,\pm]}=\chi_{1}(\gamma_{1})f^{[\pm,\pm]}\chi_{2}(\gamma_{2}).$
@@
Note that in order to pin $\mathfrak{m}$ from Theorem A down to zero it is enough to assume, in addition, that $U_{m}$ define moments of a positive distribution which is usually known a priory. Moreover, it is shown in [ChelkakDeiman] that in the case of double connected domains it is enough to know that
@@ 1.2. Main result: bordered Riemann surfaces
The next natural step after Theorem A is to consider the dimer model on a bordered Riemann surface. The obstacle for applying discrete complex analysis tools in this setup is that they require the dimer graph to satisfy geometric constrains specific to the Euclidean metric. The way around this is to consider Riemann surfaces equipped with a locally flat metric with conical singularities and the dimers graphs adapted to this metric. Dimer model on such graphs has been studied in the recent work [basok2023dimers] of the author as a part of a bigger research program initiated in [BerestyckiLaslierRayI, BerestyckiLaslierRayII]. The combination of these works provides a comprehensive description of the asymptotics of the dimer model sampled on Temperleyan graphs on Riemann surfaces (including multiply connected domains).
In general, if one applies Kenyon’s method in the aforementioned setup, one ends up with a collection of functions $f^{[\pm,\pm]}$ which are multivalued and, furthermore, may have singularities of the form $z^{\theta}$ at conical singularities of the metric. The role of $dz$ in the correlation formulae (1.2) is played by a multivalued holomorphic differential $\omega$ defined so that $|\omega|^{2}$ is the locally flat singular metric on the surface. Kenyon’s method produces harmonic differentials
@@
where $\omega^{[+]}=\omega$ and $\omega^{[-]}=\bar{\omega}$. These differentials correspond to height correlations evaluated with respect to a sign-indefinite measure. The reason for these signs to occur lies in the way how Kasteleyn theorem (and hence the combinatorial identities (1.1)) changes with the change of the topology, see [Cimasoni] and references therein.
@@
where $\phi$ is the Gaussian free field with zero boundary conditions normalized as before, while $\psi$ is a random real-valued harmonic differential on $\Sigma_{0}$ satisfying the following properties:
@@
Fix an arbitrary collection of points $x_{1},\dots,x_{N}\in\Sigma_{0}$ and define $\Sigma_{0}^{\prime}=\Sigma_{0}\smallsetminus\{x_{1},\dots,x_{N}\}$. Assume that the functions $f^{[\pm,\pm]}$ and the holomorphic differential $\omega$ used to define $U_{m}$ in (1.9) are multivalued on $\Sigma_{0}^{\prime}\times\Sigma_{0}^{\prime}\smallsetminus\mathrm{diag}$ and satisfy the following properties:
@@ 1.3. Discussion of the proof
The proof of Theorems AB is based on three ingredients. The first is a formula analogous to (1.2) where on the left-hand side $\operatorname{{\mathbb{E}}}$ is replaced by $\operatorname{\mathsf{E}}_{\mathfrak{m}}$ with unknown $\mathfrak{m}$, the field $h$ is a compactified free field with some shift, and on the right-hand side $f^{[\pm,\pm]}$ is replaced by entries of the Cauchy kernel of a perturbed Cauchy–Riemann operator acting on sections of a spin line bundle on the Schottky double. We derive this formula from a variation identity for the observable $F_{\mathfrak{m}}(v)=\operatorname{\mathsf{E}}_{\mathfrak{m}}\exp\left(i\int v\wedge dh\right)$ derived in [basok2023dimers], see Theorem 3.5 and Corollary 3.11. In the case of a planar multiply connected domain such a formula can also be derived from [berggren2025gaussian, Theorem 4.2].
The second ingredient is a square root trick that appeared in [ChelkakDeiman], see Lemma 4.1. This trick allows to derive a duality between (yet unknown) boundary conditions of $f^{[\pm,\pm]}(p,q)$ in the first and in the second variables respectively. This in its turn allows to interpret these functions as a Cauchy kernel of a certain (also unknown) line bundle on the Schottky double.
The last ingredient is an abstract argument asserting that a line bundle on a Riemann surface has a Cauchy kernel if and only if it is isomorphic to a tensor product of a spin line bundle and a flat line bundle. This implies that there is a gauge relating $f^{[\pm,\pm]}$ with a Cauchy kernel from the first ingredient, thus identifying the correlations. The point in the Jacobian parametrizing the flat line bundle determines the shift of the compactified free field and the choice of $\mathfrak{m}$. We demonstrate how this works on practice via a number of examples in Section 5.
@@ 1.4. Organization of the paper
The methods that we use require a number of basic facts from the theory of Riemann surface. In hope of making the paper as accessible and self-contained as possible we have included preliminary Section 2 where most of these facts are set out. Seasoned readers can jump straight to Section 3, which contains some preliminary results about compactified free field obtained in [basok2023dimers] and their immediate corollaries. Section 4 contains the proof of the main result. Let us emphasize that the proof, although looking rather abstract on the first glance, allows to compute the parameters from Theorem A and B in many particular cases. This is demonstrated in Section 5 dedicated to examples.

### v2

@@ Abstract.
In his seminal paper [22] published in 2000 Kenyon developed a method to study the height function of the planar dimer model via discrete complex analysis tools. The core of this method is a set of identities representing height correlations through the inverse Kasteleyn operator. In a general setup, such as considered in [11, 10], scaling limits of these identities produce a set of correlation functions written in terms of a Dirac Green’s kernel with unknown boundary conditions. It was proven in [11] that, in a simply connected domain, these correlation functions always coincide with correlation functions of the Gaussian free field given that they satisfy some natural a priori assumptions. This was generalized to doubly connected domains in the recent work [9], where correlations are shown to be the correlations of a sum of Gaussian free field and a discrete Gaussian component. We generalize this result further to arbitrary bordered Riemann surfaces.
@@
A University of Helsinki, Department of Mathematics and Statistics, Helsinki, Finland
@@ 1. Introduction
In his pioneering works [22, 23] Kenyon considered the dimer model sampled in a sequence of Temperleyan polygons $\Omega_{k}\subset\frac{1}{k}{\mathbb{Z}}^{2}$ that approximate a planar domain $\Omega\subset{\mathbb{C}}$ and proved the following two results. Assuming simply connectedness of $\Omega$, the centered dimer height functions converge to the (properly scaled) Gaussian free field with Dirichlet boundary conditions [23]. For multiply connected $\Omega$ the height gaps between boundary components of $\Omega_{k}$ converge to unknown conformally invariant random variables [22].
@@
Once one is able to make this line of arguments rigorous, one still has to identify the right-hand side of (1.2). This can be done, for example, by identifying the boundary conditions of the Dirac operator inherited by the combinatorics of discrete domains along their boundaries. For instance, Temperleyan combinatorics considered in [22, 23] imposes Dirichlet boundary conditions. These boundary conditions make $f^{[\pm,\pm]}$ conformally covariant, which implies the conformal invariance of the correlations (1.2) and allows one to identify $\overline{h}$ with the Gaussian free field in the case of a simply connected $\Omega$. As shown in a recent work [25], if $\Omega$ is multiply connected, these boundary conditions also allow one to express $f^{[\pm,\pm]}$ explicitly via the classical Szegö kernel on the Schottky double of $\Omega$ [14, Chapter VI]. This allows to fully determine the field $\overline{h}$ and, in particular, to identify the distribution of boundary height gaps considered by Kenyon in [22].
Other examples of explicit boundary conditions were studied in [27, 28]. In general, however, the boundary conditions of the dimer models in question may be quite nasty and do not produce conformally covariant scaling limits $f^{[\pm,\pm]}$. Moreover, even in the case of subdomains of regular lattices, if the boundary combinatorics of discrete domains is chosen arbitrary, there is no guarantee for the kernels $K^{-1}$ to converge or even to be bounded. This reflects the well-known prediction make by Kenyon and Okounkov [19] and asserting that the limit of the centered dimer height functions $\overline{h}$ becomes conformally invariant only after a proper choice of the conformal structure determined by the geometry of the so-called “limit shape”, i.e. the scaling limit of $\operatorname{{\mathbb{E}}}h$ [1, 20]. If the change of the conformal structure is non-trivial, one can expect the entries of $K^{-1}$ to blow up as the formulae (1.2) with the Green’s kernel being holomorphic in the initial complex structure would lead to a contradiction. Such a behaviour of $K^{-1}$ is exhibited, for example, in the case of the dimer model sampled on the Aztec diamond [6].
Having said all that, the aforementioned approach leading to formulae (1.2) can still in principle be adapted even when the change of the conformal structure is non-trivial by employing the theory of t-embeddings [11, 10, 21]. The idea is to determine a gauge of the Kasteleyn operator $K$ that will make the kernel $K^{-1}$ bounded and that will carry the geometric information needed to reconstruct the conformal structure. As suggested by [10], a natural candidate for such a gauge is a Coulomb gauge determining a perfect t-embedding of the dimer graph. The results of [10] show that, after such a gauge was applied, the resulting kernel $K^{-1}$ can be bounded in terms of the regularity of the embedding only. Regularity theory developed in [11] and results of [10] imply that subsequential scaling limits of the kernels $K^{-1}$ exist and can be identified with Dirac Green’s kernels in the conformal structure of the limiting t-surface (the limit of the lifts of perfect t-embeddings to the Lorentz space ${\mathbb{R}}^{2,2}$) with unknown boundary conditions, given that the limiting t-surface is Lorentz minimal. As discussed in  [10, Section 4.2], the latter is expected to be Lorentz minimal as soon as the dimer height function is expected to have Gaussian fluctuations in the limit. Basing on these constructions and replacing the initial $K^{-1}$ in (1.1) with the one obtained after applying the gauge, we can derive formulas (1.2) for the limits of height correlations with $f^{[\pm,\pm]}$ being components of a Dirac Green’s kernel on a Lorentz minimal surface in the Lorentz space ${\mathbb{R}}^{2,2}$.
Perfect t-embeddings have been shown to exist and exhibit the expected properties in a number of particular cases (see [12, 6, 8, 5, 18] and [21, Section 3.3]), yet the problem of their existence for general planar bipartite graphs with positive weights remains widely open. Nevertheless, the discussion above significantly strengthens the motivation to study the structure that formulas (1.2) impose on the correlations of the limiting field $\overline{h}$. While the functions $f^{[\pm,\pm]}$ that appear in these formulas may not be conformally covariant and sometimes depend on the particular subsequence along which the kernels $K^{-1}$ converge, the limiting field $\overline{h}$ is always expected to belong to a universal family of conformally invariant random generalized functions corresponding to compactified free fields in $\Omega$. This poses the problem of determining which properties of $f^{[\pm,\pm]}$ are sufficient to guarantee the conformal invariance of the right-hand side of (1.2) and, moreover, to identify the field $\overline{h}$ with a canonical object such as compactified free field.
@@
In the case of simply connected domains a compelling answer to this question is given by [11, Lemma 7.3] that asserts the following. Denote the right-hand side of (1.2) by $U_{m}$. By definition $U_{m}$ is a harmonic differential in each of the variables. Assume that $U_{2}$ and $U_{3}$ have Dirichlet boundary conditions (which one expects as $\overline{h}$ is supposed to vanish outside of the domain) and correct singularities in the bulk: that is, $U_{2}$ has a quadratic pole with a prescribed residue along the diagonal and $U_{3}$ is regular. Then there exists a gauge relating $f^{[\pm,\pm]}$ to the components of the Dirac Green’s kernel satisfying “spinor-type” boundary conditions, and, in particular, $U_{m}$’s coincide with multi-point correlations of the gradient of the Gaussian free field.
@@
When $\Omega$ is multiply connected with boundary components $B_{0},B_{1},\dots,B_{n},\ n\geq 1,$ the well-known heuristics [16, Lecture 24] suggest that the limit field $\overline{h}$ should be of the form
@@
Thus, generalizing [11, Lemma 7.3] to multiply connected domains requires to determine the distribution of the discrete component $c$ additionally to the GFF component $\phi$. In the case of a doubly connected domain such generalization was very recently obtained in [9] using properties of the Weierstrass $\wp$-function. This, in particular, allowed the authors of [9] to study the dimer model on a cylinder with Temperleyan conditions of different ‘colors’ on two boundaries, in which case the Dirac Green’s kernel $f^{[\pm,\pm]}$ is not conformally covariant (see Section 5.2).
The goal of this work is to generalize this approach further: from simply and doubly connected setups to arbitrary multiply connected domains further to bordered Riemann surfaces. In particular, we show that the algebraic structure of Kenyon’s identities (1.2) allows one to identify the form of the joint distribution of dimer height gaps between boundary components and thus provide a strong supporting evidence for [16, Conjecture 24.1].
@@ 1.1. Main result: multiply connected domains
We first formulate our theorem in the particular case of multiply connected domains. We need several definitions. Let $U_{m}$ denote the right-hand side of (1.2) as before. Let us say that $h=\phi+\sum_{j=1}^{n}c_{j}\mathrm{hm}_{j}$ is a compactified free field with the shift $c_{0}\in{\mathbb{R}}^{n}$ if $\phi$ and $c$ are as described above. Recall that the right-hand side of (1.3) is $h-\operatorname{{\mathbb{E}}}h$. Given an integer vector $\mathfrak{m}\in{\mathbb{Z}}^{n}$ define
@@
For various applications it is also convenient to allow functions $f^{[\pm,\pm]}$ to have multiplicative monodromies. (For instance, this includes the setup of [9] in which case $f^{[\pm,\pm]}$ have monodromies 1-1 in each of the variables.) Let us say that $f^{[\pm,\pm]}(z_{1},z_{2})$ are multivalued with multiplicative monodromy if $f^{[\pm,\pm]}(z_{1},z_{2})$ are functions defined for $(z_{1},z_{2})$ from the universal cover of $\Omega\times\Omega$, and there exist characters $\chi_{1,2}:\pi_{1}(\Omega)\to{\mathbb{T}}$ such that for every deck transformation $(\gamma_{1},\gamma_{2})\in\pi_{1}(\Omega)\times\pi_{1}(\Omega)$
(1.6) $(\gamma_{1},\gamma_{2})^{\ast}f^{[\pm,\pm]}=\chi_{1}(\gamma_{1})\chi_{2}(\gamma_{2})f^{[\pm,\pm]}.$
@@
Note that in order to pin $\mathfrak{m}$ from Theorem A down to zero it is enough to assume, in addition, that $U_{m}$ define moments of a positive distribution which is usually known a priory. Moreover, it is shown in [9] that in the case of double connected domains it is enough to know that
@@ 1.2. Main result: bordered Riemann surfaces
The next natural step after Theorem A is to consider the dimer model on a bordered Riemann surface. The obstacle for applying discrete complex analysis tools in this setup is that they require the dimer graphs to be adapted to the Euclidean metric on the plane. The way around this is to consider Riemann surfaces equipped with a locally flat metric with conical singularities. Dimer model in such a setup has been studied in the recent work [2] of the author as a part of a bigger research program initiated in [3, 4]. The combination of these works provides a comprehensive description of the asymptotics of the dimer model sampled on Temperleyan graphs on Riemann surfaces (including multiply connected domains).
In general, if one applies Kenyon’s method in the aforementioned setting, one ends up with a collection of functions $f^{[\pm,\pm]}$ which are multivalued and, furthermore, may have singularities of the form $z^{\theta}$ at conical singularities of the metric. The role of $dz$ in the correlation formulae (1.2) is played by a multivalued holomorphic differential $\omega$ defined so that $|\omega|^{2}$ is the locally flat singular metric on the surface. Kenyon’s method produces harmonic differentials
@@
where $\omega^{[+]}=\omega$ and $\omega^{[-]}=\bar{\omega}$. These differentials correspond to height correlations evaluated with respect to a sign-indefinite measure. The reason for the signs in the measure to occur lies in the way how the Kasteleyn theorem (and hence the combinatorial identities (1.1)) changes when the topology is non-trivial, see [13] and references therein.
@@
where $\phi$ is the Gaussian free field with zero boundary conditions normalized as before and $\psi$ is a random real-valued harmonic differential on $\Sigma_{0}$ satisfying the following properties:
@@
Fix an arbitrary collection of points $x_{1},\dots,x_{N}\in\Sigma_{0}$ and define $\Sigma_{0}^{\prime}=\Sigma_{0}\smallsetminus\{x_{1},\dots,x_{N}\}$. Assume that the functions $f^{[\pm,\pm]}$ and the holomorphic differential $\omega$ used to define $U_{m}$ in (1.9) are multivalued on $\Sigma_{0}^{\prime}\times\Sigma_{0}^{\prime}\smallsetminus\mathrm{diag}$ and $\Sigma_{0}^{\prime}$ respectively and satisfy the following properties:
@@ 1.3. Discussion of the proof
The proof of Theorems A and B is based on three ingredients. The first is a formula analogous to (1.2) where on the left-hand side $\operatorname{{\mathbb{E}}}$ is replaced by $\operatorname{\mathsf{E}}_{\mathfrak{m}}$ with unknown $\mathfrak{m}$, the field $h$ is a compactified free field with some shift, and on the right-hand side $f^{[\pm,\pm]}$ is replaced by entries of the Cauchy kernel of a perturbed Cauchy–Riemann operator acting on sections of a spin line bundle on the Schottky double of the Riemann surface. We derive this formula from a variation identity for the observable $F_{\mathfrak{m}}(v)=\operatorname{\mathsf{E}}_{\mathfrak{m}}\exp\left(i\int v\wedge dh\right)$ derived in [2], see Theorem 3.5 and Corollary 3.11. In the case of a planar multiply connected domain such a formula can also be derived from [7, Theorem 4.2].
The second ingredient is a square root trick that appeared in [9, Theorem 3.8], see Lemma 4.1. This trick allows to derive a duality between (yet unknown) boundary conditions of $f^{[\pm,\pm]}(p,q)$ in the first and in the second variables respectively. This in its turn allows to interpret these functions as a Cauchy kernel of a certain (also unknown) line bundle on the Schottky double.
The last ingredient is an abstract argument asserting that a line bundle on a Riemann surface has a Cauchy kernel if and only if it is isomorphic to a tensor product of a spin line bundle and a flat line bundle. This implies that there is a gauge relating $f^{[\pm,\pm]}$ with a Cauchy kernel from the first ingredient, thus identifying the correlations. The point in the Jacobian parametrizing the flat line bundle determines the shift of the compactified free field and the choice of $\mathfrak{m}$. We demonstrate how this works in practice via a number of examples in Section 5.
To conclude the discussion, let us note that Theorem B holds also for non-bordered surfaces. In fact, it is much easier in this case as one does not need to deal with boundary conditions of $f^{[\pm,\pm]}$, thus the second (the most technical) step can be omitted and the theorem becomes a simple corollary of the calculations made in Section 3.4.
@@ 1.4. Organization of the paper
The methods that we use require a number of basic facts from the theory of Riemann surface. In hope of making the paper as self-contained as possible we have included preliminary Section 2 where most of these facts are set out. Seasoned readers can jump straight to Section 3, which contains some preliminary results about compactified free field obtained in [2] and their immediate corollaries. Section 4 contains the proof of the main result. Let us emphasize that the proof, although looking rather abstract on the first glance, allows to compute the parameters from Theorem A and B in many particular cases. This is demonstrated in Section 5 dedicated to examples.